# Evaluation Should Not Ignore Variation: On the Impact of Reference Set Choice on Summarization Metrics

### v1

@@ Abstract
Human language production exhibits remarkable richness and variation, reflecting diverse communication styles and intents. However, this variation is often overlooked in summarization evaluation. While having multiple reference summaries is known to improve correlation with human judgments, the impact of using different reference sets on reference-based metrics has not been systematically investigated. This work examines the sensitivity of widely used reference-based metrics in relation to the choice of reference sets, analyzing three diverse multi-reference summarization datasets: SummEval, GUMSum, and DUC2004. We demonstrate that many popular metrics exhibit significant instability. This instability is particularly concerning for n-gram-based metrics like ROUGE, where model rankings vary depending on the reference sets, undermining the reliability of model comparisons. We also collect human judgments on LLM outputs for genre-diverse data and examine their correlation with metrics to supplement existing findings beyond newswire summaries, finding weak-to-no correlation. Taken together, we recommend incorporating reference set variation into summarization evaluation to enhance consistency alongside correlation with human judgments, especially when evaluating LLMs.
Evaluation Should Not Ignore Variation:
On the Impact of Reference Set Choice on Summarization Metrics
@@ 1 Introduction
Figure 1: Human-written summaries are diverse. Using a human-written reference instead of another makes evaluation metrics fluctuate, and affects model ranking.
they remain widely popular due to their simplicity, low compute requirements, relative ease of adaptation to different languages, and reproducibility. In 2024, out of $21$ ACL papers mentioning “summarization” in their title, $19$ ($90$%) include a reference-based metric in their evaluation, with ROUGE Lin (2004) being the most common ($71$%), followed by BERTScore (Zhang et al. 2020, $52$%).
The assumption behind the use of reference-based metrics is that systems outputs that are more similar to the reference(s) are better, due to their “human-likeness” Gehrmann et al. (2023). However, the significant variation in human-written summaries implies that evaluating system outputs against a single or limited set of references has inherent drawback. Previous research has extensively looked at correlation between metrics and human judgments in summarization, further exploring the use of multiple references to improve such correlation Lin (2004); Belz and Reiter (2006); Fabbri et al. (2021); Tang et al. (2024). However, a much less studied question is the extent to which automatic metrics are sensitive to the choice of human-written reference summaries, as shown in Figure 1. In other words, are these metrics stable across different plausible gold-standard references? If metric scores vary significantly with the selected reference(s), this variation calls into question the reliability of many evaluation practices in the field.
@@
    We investigate how different reference sets affect system rankings. Our results show that system rankings based on n-gram-matching metrics (e.g., ROUGE) strongly depend on the choice of the reference(s), undermining the reliability of model comparisons. However, rankings based on more semantically-oriented metrics exhibit greater stability.
@@
    We examine the robustness of widely-used reference-based metrics at the instance and dataset level. Our analysis reveals that the variation in scores introduced by the choice of reference on a dataset often exceeds the variation observed in state-of-the-art (SOTA) summarization models.
@@
    We collect new human judgment scores on Large Language Model (LLM) outputs for the genre-diverse GUMSum Liu and Zeldes (2023) dataset. We use these data to reassess the correlation between automatic metrics and human judgments, complementing earlier SummEval evaluation (Fabbri et al., 2021) limited to pre-LLM models and newswire data. We find that correlations tend to increase with the number of references, and that the metric with the highest correlation varies depending on the evaluation dimension and the number of references.
Our analysis reveals that few metrics tend to show reasonable correlation with human judgments and robustness to the reference sets, especially when scoring LLM outputs.
@@ 2 Related Work
Recent advances in Natural Language Generation (NLG) have significantly enhanced the development of automatic summarization systems. However, their evaluation remains an open problem Celikyilmaz et al. (2021); Goyal et al. (2023). Summarization evaluation metrics are broadly categorized into reference-based and reference-free Lloret et al. (2018). Reference-based metrics compare system outputs to human-written reference summaries, relying on methods such as n-gram overlap Lin (2004); Papineni et al. (2002), embedding similarity Ng and Abrecht (2015); Zhao et al. (2019); Zhang et al. (2020), or model-based evaluation techniques Peyrard et al. (2017); Scialom et al. (2019); Yuan et al. (2021). In contrast, reference-free summarization metrics do not assume a gold standard Yuan et al. (2021); Vasilyev et al. (2020); Gao et al. (2020). Lastly, growing research starts to leverage LLMs as evaluators, with or without references Liu et al. (2023); Song et al. (2024); Li et al. (2024).
@@
Meta-evaluation of summarization metrics typically focuses on the extent to which they can be used as a proxy for human evaluation. Reiter and Belz (2009) examined the validity of automatic scores for NLG tasks, while Rankel et al. (2013) focused on ROUGE and its correlation with human judgments. Peyrard (2019) showed that metrics with reasonable correlation on lower-quality outputs tend to diverge when output quality increases. Caglayan et al. (2020) demonstrated various idiosyncrasies of automatic evaluation metrics, noting that high correlation with human judgments is not sufficient to characterize their reliability. Fabbri et al. (2021) performed a large-scale meta-evaluation of summarization metrics, and found that most metrics have low correlation with human judgments on coherence, while relevance is weakly or moderately correlated.
@@ 3 Experimental Setup
To quantify the impact of different human-written references on the scores of automatic metrics, we exam multiple elements. For datasets, we use SummEval Fabbri et al. (2021), GUMSum Liu and Zeldes (2023), and DUC2004 Dang and Croft (2004), which contain multiple human-written summaries (§3.1), to assess how different reference summaries affect metric performances. Next, to assess summarization models, we use the existing outputs provided by Fabbri et al. (2021) for SummEval. As these outputs predate LLMs, we additionally collect outputs using LLMs (§3.2) for all three datasets. Lastly, to compute the correlations with humans, we use the human judgments available in SummEval and gather new human ratings for GUMSum on both human and LLM-generated summaries (§3.3). We prioritize GUMSum over DUC2004, as it includes multiple genres beyond news data. We comply with the license of the existing datasets. For newly collected model outputs and human judgments, we follow the license of the corresponding underlying datasets. Our metric selection is outlined in §3.4.
@@ 3.1 Human-written Summaries
SummEval Fabbri et al. (2021) is built on top of CNN/DM Hermann et al. (2015); Nallapati et al. (2016), containing news articles and human-written highlights. Authors selected $100$ instances from the test set; for these, in addition to the highlight-based summary in CNN/DM, ten references were crowd-sourced Kryscinski et al. (2019).
GUMSum Liu and Zeldes (2023) contains summaries created following general and genre-specific guidelines [^1] https://wiki.gucorpling.org/gum/summarization to function as a substitute for the source Nenkova and McKeown (2011). We focus on the $48$ documents in the dev and test sets, which contain five human-written summaries each Lin and Zeldes (2025), evenly distributed across $12$ genres.
DUC2004 Task1 Dang and Croft (2004) consists of $489$ news documents, most with four references. The guidelines allow the summaries to be in the form of short sentences or lists of keywords. [^2] https://duc.nist.gov/duc2004/tasks.html DUC2004 references are thus extremely concise (only up to $75$ characters). The dataset has played a significant role in summarization research and was part of the annual TREC conference evaluation.
@@ 3.2 Model Outputs
Fabbri et al. (2021) collected model outputs for SummEval from $24$ extractive and abstractive summarization systems, which were SOTA between 2017 and 2019. We focus on the $16$ models for which they provided human judgments.
For all datasets, we also include summaries generated by contemporary LLMs. This is crucial given that prior studies have demonstrated that evaluation metrics often show lower correlation with high-quality outputs Peyrard (2019); Alva-Manchego et al. (2021). We see a similar pattern for LLMs (§4.4). For consistency purposes, we follow Lin and Zeldes (2025) and use Llama3-3B-Instruct Hermann et al. (2015), Qwen-2.5-7B-Instruct Qwen et al. (2025), Claude-3.5 Anthropic (2024), and GPT-4o OpenAI (2024). For each LLM, we generate a single summary. This way, we emphasize LLM variety over multiple generations. Details on the generation parameters and prompts etc. are reported in Appendix A.
@@ 3.3 Human Judgments
SummEval (Fabbri et al., 2021) contains expert judgments that assess summaries based on four criteria: coherence, consistency, fluency, and relevance, using a Likert scale of $1$-$5$ Likert (1932).
To measure how well automatic metrics align with human judgments in different genres, and to study whether findings on pre-LLM models align with those on LLM outputs, we conduct a human evaluation on the $48$ GUMSum documents. We recruited three Master’s students in Computational Linguistics and instructed them to evaluate four LLM outputs (§3.2) and five human references, following Fabbri et al. (2021)’s criteria.
@@
Table 2: Human judgments on GUMSum for system-generated versus human-written summaries. best and worst indicate the percentage of evaluators who voted the summarizer as best/worst averaged over documents.
LLM-generated and human-written summaries were anonymized and shuffled. We also asked the evaluators to pick one best and one worst summary for each document. Table 2 reports the results. We note that overall, Claude scored the best. GPT-4o gets the highest consistency but the lowest coherence and relevance, and thus the least picked LLM output. Interestingly, LLM outputs typically receive higher scores than human-written references.
@@ 3.4 Evaluation Metrics
We examine the following reference-based metrics chosen due to their widespread use. All metric ranges fall in 00-$100$. Appendix B provides details.
ROUGE Lin (2004) is the most popular metric for summarization. ROUGE-N computes n-gram overlap between a hypothesis and the references. ROUGE-L leverages the longest common subsequence, accounting for the word order. When evaluating with multiple references, ROUGE considers either the maximum or the mean of the n-gram overlap (ROUGEmax and ROUGEavg). We report the F1-score.
BLEU Papineni et al. (2002) is an n-gram overlap metric primarily used to assess translations. It is precision-based and incorporates a brevity penalty. When multiple references are provided, the n-gram count is clipped at the maximum count of n-grams in a single reference and the length of the reference closest in size to the hypothesis is considered.
@@
BERTScore Zhang et al. (2020) leverages pre-trained contextual embeddings and considers the cosine similarity between the embeddings of the hypothesis and the reference tokens. With multiple references, the final score is the maximum among the individual scores. We report the F1 score.
BLEURT (Sellam et al., 2020) is a model-based metric that leverages a pre-trained BERT fine-tuned on human judgments. The metric is not designed to handle multiple references; thus, we compute individual scores for each reference and consider the maximum value.
4 Reference Variability and Metrics Robustness
@@ 4.1 Human-written Summaries are Diverse
Human-written summaries show substantial diversity. We assess the variability in the multi-reference datasets following Giulianelli et al. (2023). For each pair of human-written summaries for the same instance in the three datasets, we report the lexical similarity (the overlapping distinct n-grams between two strings), the syntactic similarity (the overlap of part-of-speech tag n-grams), and the semantic similarity (the cosine and euclidean similarity between the embeddings of the two strings).
Figure 2 shows these variations. At the dataset level, DUC and SummEval show the lowest similarity among human-written summaries across all dimensions. For GUMSum, summaries are more similar to each other. We hypothesize that this is likely due to the constrained annotation guidelines. [^3] https://wiki.gucorpling.org/gum/summarization It is also worth noting that the similarities revealed here are between different human-written summaries for a given instance as opposed to summaries across genres, for which we still expect significant variations, as demonstrated by Liu and Zeldes (2023). Overall, summaries tend to be similar at the syntactic level, less so at the semantic and lexical level. We also observe that LLM outputs show lower diversity (Appendix C), consistently with previous work Giulianelli et al. (2023).
@@ 4.2 Automatic Metrics Fluctuate Substantially at the Instance Level
Given the diversity in human-written summaries, we aim to quantify metric fluctuations at the instance level when using a different set of human-written references. For an automatic evaluation metric $M$ and a set of human-written references $R=\{r_{1},r_{2},\dots,r_{N}\}$, we compute $M(r_{i},R-\{r_{i}\})$. In other words, for each document, we score each human-written summary using all the others as the reference set. Figure 3 exemplifies the observed variability at the instance level measured by ROUGE-Lavg on the three datasets. For SummEval, we also highlight the scraped original reference in the CNN/DM dataset with a cross. The quality of these scraped references versus the ten crowd-sourced ones is discussed further in Appendix D.
Figure 3: Instance-level variation for ROUGE-Lavg. For every document (shown first $30$, one per line), we plot the score for every human-written reference against all other references (using the same color per source to aid interpretation). The original CNN/DM reference in SummEval is marked by a cross.
@@
Human-written references scores vary widely.
Figure 4 summarizes the instance-level variability of the individual scores (in Figure 3) for all evaluation metrics on SummEval (corresponding figures for GUMSum and DUC are in Appendix E). For each metric, we compute the range (i.e., the difference between the maximum and the minimum score) when scoring human-written references against all the others ($M(r_{i},R-\{r_{i}\})$). Figure 4 shows the histogram of such ranges. Note that the ranges of variation observed within human-written references are on average very high.
Understanding the magnitude of such a range might not be obvious. For instance, an increase of $10$ points of BERTScore (which typically returns scores compressed in the high range of the scale) might indicate a much larger improvement in performance than an increase of $10$ points of ROUGE-1.444The relative dynamics between metrics have been studied by Kocmi et al. (2024) in machine translation. To contextualize the magnitude of variation for each metric, we also report the range of performance of summarization systems. Thus, for a model $S$, given its output $o_{i}$ for instance $i$, we score it through $M(o_{i},R)$. Although these values are not directly comparable and should be interpreted with caution due to the use of different reference sets, they help contextualize the magnitude of the results and its potential impact on evaluation. For example, ROUGE-1max assigned to human-written references vary by about $35$ points on average (the green dashed line in Figure 4). In contrast, the mean range is less than $20$ points across all model outputs (orange line), and much lower for LLM-generated outputs (blue line). These findings highlight the significance of the observed variability and suggest that the ranking of summarization models is highly sensitive to the reference set.

### v2

# References Matter: Investigating the Impact of Reference Set Variation on Summarization Evaluation

@@ Abstract
Human language production exhibits remarkable richness and variation, reflecting diverse communication styles and intents. However, this variation is often overlooked in summarization evaluation. While having multiple reference summaries is known to improve correlation with human judgments, the impact of the reference set on reference-based metrics has not been systematically investigated. This work examines the sensitivity of widely used reference-based metrics in relation to the choice of reference sets, analyzing three diverse multi-reference summarization datasets: SummEval, GUMSum, and DUC2004. We demonstrate that many popular metrics exhibit significant instability. This instability is particularly concerning for n-gram-based metrics like ROUGE, where model rankings vary depending on the reference sets, undermining the reliability of model comparisons. We also collect human judgments on LLM outputs for genre-diverse data and examine their correlation with metrics to supplement existing findings beyond newswire summaries, finding weak-to-no correlation. Taken together, we recommend incorporating reference set variation into summarization evaluation to enhance consistency alongside correlation with human judgments, especially when evaluating LLMs.
References Matter: Investigating the Impact of Reference Set Variation
on Summarization Evaluation
@@ 1 Introduction
Figure 1: Human-written summaries are diverse. Using a human-written reference over another makes evaluation metrics fluctuate, and affects model ranking.
they remain widely popular due to their simplicity, low compute requirements, relative ease of adaptation to different languages, and reproducibility.
The assumption behind the use of reference-based metrics is that system outputs that are more similar to the reference(s) are better, due to their “human-likeness” Gehrmann et al. (2023). However, the significant variation in human-written summaries implies that evaluating system outputs against a single or limited set of references has inherent drawbacks. Previous research has extensively looked at correlation between metrics and human judgments in summarization Forde et al. (2024), further exploring the use of multiple references to improve such correlations Lin (2004); Belz and Reiter (2006); Fabbri et al. (2021); Tang et al. (2024) as well as the interpretability and efficiency aspects of such automatic metrics Liu et al. (2023b). However, a much less studied question is the extent to which automatic metrics are sensitive to the choice of human-written reference summaries, as shown in Figure 1. In other words, are these metrics stable across different plausible gold-standard references? If metric scores vary significantly with the selected reference(s), this variation calls into question the reliability of many evaluation practices in the field.
@@
    We investigate how different reference sets affect system rankings. We show that system rankings based on n-gram-matching metrics (e.g., ROUGE) strongly depend on the choice of the reference(s), undermining the reliability of model comparisons. However, rankings based on more semantically-oriented metrics exhibit greater stability.
@@
    We examine the robustness of widely-used reference-based metrics for summarization at the instance and dataset level. Our analysis reveals that the variation in scores introduced by the choice of reference on a dataset often exceeds the variation observed in state-of-the-art (SOTA) models.
@@
    We collect new human judgment scores on Large Language Model (LLM) outputs for the genre-diverse GUMSum Liu and Zeldes (2023) dataset. We use these data to reassess the correlation between automatic metrics and human judgments, complementing earlier SummEval evaluations (Fabbri et al., 2021), which were limited to pre-LLM models and newswire data. We find that correlations tend to increase with the number of references, and that the metric with the highest correlation varies depending on the evaluation dimension and the number of references.
Our analysis reveals that few metrics tend to show both reasonable correlation with human judgments and robustness to the reference sets, especially when scoring LLM outputs.
@@ 2 Related Work
Recent advances in Natural Language Generation (NLG) have significantly enhanced the development of summarization systems. However, their evaluation remains an open problem Celikyilmaz et al. (2021); Goyal et al. (2023). Summarization evaluation metrics are broadly categorized into reference-based and reference-free Lloret et al. (2018). Reference-based metrics compare system outputs to human-written reference summaries, relying on methods such as n-gram overlap Lin (2004); Papineni et al. (2002), embedding similarity Ng and Abrecht (2015); Zhao et al. (2019); Zhang et al. (2020), or model-based techniques Peyrard et al. (2017); Scialom et al. (2019); Yuan et al. (2021). In contrast, reference-free summarization metrics do not assume a gold standard Yuan et al. (2021); Vasilyev et al. (2020); Gao et al. (2020). Lastly, growing research leverages LLMs as evaluators, with or without references Liu et al. (2023a); Song et al. (2024); Li et al. (2024); Gigant et al. (2024).
@@
Meta-evaluation of summarization metrics typically focuses on the extent to which they can be used as a proxy for human evaluation. Reiter and Belz (2009) examined the validity of automatic scores for NLG tasks, while Rankel et al. (2013) focused on ROUGE and its correlation with humans. Peyrard (2019) showed that metrics with reasonable correlation on lower-quality outputs tend to diverge when output quality increases. Caglayan et al. (2020) demonstrated the idiosyncrasies of automatic evaluation metrics, noting that high correlation with human judgments is not sufficient to characterize their reliability. Fabbri et al. (2021) performed a large-scale meta-evaluation of summarization metrics, and found that most metrics have low correlation with human judgments on coherence, while relevance is weakly or moderately correlated.
@@ 3 Experimental Setup
To quantify the impact of human-written references on the scores of automatic metrics, we exam multiple elements. For datasets, we use SummEval Fabbri et al. (2021), GUMSum Liu and Zeldes (2023), and DUC2004 Dang and Croft (2004), which contain multiple human-written summaries (§3.1), to assess how different reference summaries affect metric performances. Next, to assess summarization models, we use the existing outputs provided by Fabbri et al. (2021) for SummEval. As these outputs predate LLMs, we additionally collect outputs using LLMs (§3.2) for all three datasets. Lastly, to compute the correlations with humans, we use the human judgments available in SummEval and gather new human ratings for GUMSum on both human and LLM-generated summaries (§3.3). We prioritize GUMSum over DUC2004, as it includes multiple genres beyond news data. Our metric selection is outlined in §3.4. Details on data licensing and codebase are provided in Appendix A.
@@ 3.1 Human-written Summaries
SummEval Fabbri et al. (2021) is built on top of CNN/DM Hermann et al. (2015); Nallapati et al. (2016), containing news articles and human-written highlights. The original authors selected 100100100 instances from the test set and supplemented the existing highlights with ten additional reference summaries per instance, obtained via crowd-sourcing Kryscinski et al. (2019).
GUMSum Liu and Zeldes (2023) contains summaries created following general and genre-specific guidelines to function as a substitute for the source Nenkova and McKeown (2011). We focus on the 484848 documents in the dev and test sets, which contain five human-written summaries each Lin and Zeldes (2025) across 121212 genres.
DUC2004 Task1 Dang and Croft (2004) consists of 489489489 news documents, most with four references. The guidelines allow the summaries to be in the form of short sentences or lists of keywords. [^1] https://duc.nist.gov/duc2004/tasks.html DUC2004 references are thus extremely concise (only up to 757575 characters). The dataset has played a significant role in summarization research, being part of the annual TREC conference evaluation.
@@ 3.2 Model Outputs
Fabbri et al. (2021) collected model outputs for SummEval from 242424 extractive and abstractive summarization systems, which were SOTA between 2017 and 2019. We focus on the 161616 models for which they provided human judgments.
For all datasets, we also include summaries generated by contemporary LLMs. This is crucial given that prior studies demonstrated that evaluation metrics often show lower correlation with high-quality outputs Peyrard (2019); Alva-Manchego et al. (2021). Below, we report a similar pattern for LLMs (§4.4). For consistency purposes, we follow Lin and Zeldes (2025) and use Llama3-3B-Instruct Hermann et al. (2015), Qwen-2.5-7B-Instruct Qwen et al. (2025), Claude-3.5 Anthropic (2024), and GPT-4o OpenAI (2024). For each LLM, we generate a single summary. We emphasize LLM variety over multiple generations. Details on the generation parameters and prompts are reported in Appendix A.
@@ 3.3 Human Judgments
SummEval (Fabbri et al., 2021) contains expert judgments that assess summaries based on four criteria: coherence, consistency, fluency, and relevance, using a Likert scale of 111-555 Likert (1932).
To measure how well automatic metrics align with human judgments in different genres, and to study whether findings on pre-LLM models align with those on LLM outputs, we collect a new set of human judgments using the same criteria on the 484848 GUMSum documents. We hired three Master’s students in Computational Linguistics and tasked them to evaluate four LLM outputs (§3.2) and five human references, following Fabbri et al. (2021)’s criteria.
@@
Table 2: Human judgments on GUMSum for system- versus human-written summaries.
LLM-generated and human-written summaries were anonymized and shuffled. We also asked the evaluators to pick one best and one worst summary for each document. Table 2 reports the results. Claude scored the best overall. GPT-4o gets the highest consistency but the lowest coherence and relevance, and thus the least picked LLM. Interestingly, LLM outputs typically receive higher scores than human-written references, in line with existing research Zhang et al. (2024).
@@ 3.4 Evaluation Metrics
To gain insights into recent metric usage, Figure 2 summarizes the percentage of summarization papers from recent ACL/EMNLP/INLG proceedings (detailed in Appendix C). We found that reference-based metrics are still the most popular, with ROUGE in 797979% of papers, followed by BERTScore (444444%). Our preliminary search also shows that LLM-as-a-judge evaluators are mainly (i.e., in 666666% of the cases) used without references, placing them outside the scope of this paper. Below, we describe the commonly used reference-based metrics in the surveyed papers, balancing popularity and diversity. All metrics range in 0100100100. Appendix B provides details.
ROUGE Lin (2004) is the most popular summarization metric. ROUGE-N computes n-gram overlap between a hypothesis and the references. ROUGE-L leverages the longest common subsequence, accounting for the word order. With multiple references, ROUGE considers the maximum or the mean of the n-gram overlap (ROUGEmax and ROUGEavg). We report the F1-score.
BLEU Papineni et al. (2002) is an n-gram metric primarily used for translation. It is precision-based and incorporates a brevity penalty. With multiple references, the n-gram count is clipped at the maximum count of n-grams in a single reference and the length of the reference closest in size to the hypothesis is considered.
@@
BERTScore Zhang et al. (2020) leverages contextual embeddings and considers the cosine similarity between the embeddings of the hypothesis and the reference tokens. With multiple references, the final score is the maximum among the individual scores. We report the F1 score.
BLEURT (Sellam et al., 2020) is a model-based metric that leverages BERT fine-tuned on human judgments. The metric is not designed to handle multiple references; we compute scores for each reference and consider the maximum.
Figure 2: Percentage of papers about summarization that use common reference-based metrics.
4 Reference Variability and Metric Robustness
@@ 4.1 Human-written Summaries are Diverse
Human-written summaries show substantial diversity. We assess the variability in the multi-reference datasets following Giulianelli et al. (2023). For each pair of human-written summaries for the same instance, we report the lexical similarity (the overlapping distinct n-grams between two strings), the syntactic similarity (the overlap of part-of-speech tag n-grams), and the semantic similarity (the cosine and euclidean similarity between the embeddings of the two strings).
Figure 3 shows these variations. At the dataset level, DUC and SummEval show the lowest similarity among human-written summaries across all dimensions. For GUMSum, summaries are more similar to each other. We hypothesize that this is likely due to the constrained annotation guidelines. [^2] https://wiki.gucorpling.org/gum/summarization It is also worth noting that the similarities revealed here are between different human-written summaries for a given instance as opposed to summaries across genres, for which we still expect significant variations, as demonstrated by Liu and Zeldes (2023). Overall, summaries tend to be similar at the syntactic level, less so at the semantic and lexical level. We also observe that LLM outputs show lower diversity (Appendix D), consistently with previous work Giulianelli et al. (2023).
@@ 4.2 Automatic Metrics Fluctuate Substantially at the Instance Level
Given the diversity in human-written summaries, we quantify metric fluctuation at the instance level when using a different set of human-written references. For a metric $M$ and a set of human-written references $R=\{r_{1},r_{2},\dots,r_{N}\}$, we compute $M(r_{i},R-\{r_{i}\})$. Thus, for each document, we score each human-written summary using all the others as the reference set. Figure 4 exemplifies the observed instance-level variability measured by ROUGE-Lavg on the three datasets. For SummEval, we also highlight the scraped original reference in the CNN/DM dataset with a cross. The quality of these references versus the ten crowd-sourced ones is discussed further in Appendix E.
Figure 4: Instance-level variation for ROUGE-Lavg. For every document (shown first 303030, one per line), we plot the score for every human-written reference against all other references (using the same color per source to aid interpretation). The original CNN/DM reference in SummEval is marked by a cross.
@@
Human-written reference scores vary widely.
Figure 5 summarizes the instance-level variability of the individual scores (in Figure 4) for all evaluation metrics on SummEval (corresponding figures for GUMSum and DUC are in Appendix F). For each metric, we compute the min-max range when scoring human-written references against all the others ($M(r_{i},R-\{r_{i}\})$). Figure 5 shows the histogram of such ranges. The ranges of variation observed within human-written references are, on average, very high.
Understanding the magnitude of such a range might not be obvious. For instance, an increase of 101010 points of BERTScore (typically scoring in the high range of the scale) might indicate a much larger improvement in performance than an increase of 101010 points of ROUGE-1.333The relative dynamics between metrics have been studied by Kocmi et al. (2024) in machine translation. To contextualize the magnitude of variation for each metric, we also report the performance range of summarization systems. Thus, for a model $S$, given its output $o_{i}$ for instance $i$, we score it through $M(o_{i},R)$. Although these values are not directly comparable and should be interpreted with caution due to the use of different reference sets, they help contextualize the magnitude of the results and its potential impact on evaluation. For example, ROUGE-1max assigned to human-written references vary by about 353535 points on average (the green dashed line in Figure 5), while the mean range is less than 202020 points across all model outputs (orange line), and much lower for LLM outputs (blue line). Similarly, LLM summaries exhibit much less variability than human references on all metrics. These findings highlight the significance of variability and suggest that the ranking of summarization models is highly sensitive to the reference set.